OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (jhannybean):

distribute the -3 to both (x+5) and (x+1)

OpenStudy (jhannybean):

(-3)(x) -(3)(3) = (-3)(x) -(3)(1) -5 What do you get?

OpenStudy (anonymous):

is it -6

OpenStudy (jhannybean):

-3x -9 = -3x - 3 - 5 Add 3x to both sides -3x + 3x -9 = -15 Simplify -9 = - 15 No solution.

OpenStudy (anonymous):

thanks

OpenStudy (jhannybean):

np